# On Picard’s Theorem Via Nevanlinna Theory

### v6

@@ Abstract.
We study the classical Picard’s problem for complete Kähler manifolds with non-negative Ricci curvature. Based on the global Green function method, we generalize the Nevanlinna theory to complete Kähler manifolds with non-negative Ricci curvature. As a consequence, we give a positive answer to the Picard’s problem in the non-parabolic case, i.e., if the (non-compact) manifold is non-parabolic, then we show that every meromorphic function is a constant if it omits three distinct values.
@@
Nevanlinna theory; meromorphic mappings; Picard’s theorem; defect relation; Kähler manifolds2010 Mathematics Subject Classification:
32H30; 32H25; 32A22Contents
1. Introduction
1.1. Motivations and Techniques
The famous Picard’s theorem asserts that a meromorphic function on the complex Euclidean spaces will reduce to a constant, if it omits three distinct values. Historically, people have been committed to developing this theorem for a long time (see [4, 16, 18, 20, 21, 24, 31]). The study of Picard’s problem on manifolds may have originated from the celebrated work due to S. T. Yau [48] on Liouville’s theorem, which proves the Liouville property of harmonic functions on non-negatively Ricci curved manifolds. Yau’s work has caused a classical problem: does Picard’s theorem hold on complete Kähler manifolds with non-negative Ricci curvature? It is well-known that Nevanlinna theory [30] is a great development of both Liouville’s theorem and Picard’s theorem. So, a further problem is, can Nevanlinna theory extend to complete Kähler manifolds with non-negative Ricci curvature? As far as we know, neither of them has been solved.
Let be a complete non-compact Kähler manifold with non-negative Ricci curvature (see such examples in Sha-Yang [38] and Tian-Yau [43, 44]). The classical Picard’s problem states that
@@ 1.1. Motivations and Techniques
Remark. Let be the universal covering, where is the unit disc in Thanks to the Liouville’s theorem of Yau, we may have an affirmative answer to the Picard’s problem if every holomorphic mapping can lift to a holomorphic mapping by By the mapping lifting theorem, admits a lifting if and only if
for some such that Since is simply-connected, then the fundamental group is trivial. It implies that is necessarily a trivial group if such a lifting exists. However, is not trivial in most cases. Thus, we cannot lift to in general.
An early result in this direction could be dated back to 1970, S. Kobayashi [24] obtained some Picard-type theorems for holomorphic mappings between complex manifolds. In particular, he showed that
@@
Kobayashi’s result is subject to certain complex manifolds which are acted on transitively by a complex Lie group. In 1975, Goldberg-Ishihara-Petridis [21] established some Picard-type theorems on the locally flat manifolds (see N. Petridis [31] also). They showed that
@@
If is locally flat, then every holomorphic mapping of bounded dilatation must be a constant mapping.
In 2010, A. Atsuji [2] gave an affirmative answer for a class of meromorphic functions of slow growth rate by developing a Nevanlinna-type theory based on heat diffusion, Let us introduce his main result. Let be the Kähler form and Laplace-Beltrami operator of associated to respectively. One can endow with the Fubini-Study metric Let be a meromorphic function. The Kählerness of implies the Hilbert-Schmidt norm of the differential (with respect to ) can be formulated as
@@
Atsuji’s method [2] (see [12] also) is using stochastic calculus by Brownian motion, who introduced the so-called Nevanlinna-type functions and where is the time of the Brownian motion on By computing curvatures and using Itô’s formula (see [22]), he established a heat diffusion version of the Second Main Theorem which leads to the above Theorem C. Note that is stochastically complete (i.e., is conservative), since has non-negative Ricci curvature due to Grigor’yan’s criterion [17]. However, in order to make and meaningful in the sense of Nevanlinna theory, the following conditions are necessary:
@@
To do so, has to satisfy certain growth assumptions. That is why a growth condition is needed in Atsuji’s theorem. Unfortunately, the growth condition seems so strong that there are only a few meromorphic functions satisfy it.
The main purpose of the present paper is to generalize Nevanlinna theory to complete Kähler manifolds with non-negative Ricci curvature. It is known that Nevanlinna theory in several complex variables (see [28, 33]) studies the value distribution of meromorphic mappings between complex spaces, which has been developed for a long time. During this period, fruitful results have been achieved. For example, the reader may refer to L. Ahlfors [5], H. Cartan [7], Carlson-Griffths-King [8, 19], J. Noguchi [27, 28], E. Nochka [29], M. Ru [32, 33], B. Shiffman [34], B. Shabat [35], F. Sakai [36, 37], W. Stoll [41, 42], P. Vojta [40] and H. Wu [47], etc.; and also refer to [1, 2, 3, 12, 13, 14, 15]. To our knowledge, almost all results in Nevanlinna theory are built on complex manifolds with a complete Kähler metric of non-positive sectional curvature, while we seem to know little if the domain manifolds are not non-positively curved, since it is difficult to estimate the local Green function for a general bounded domain, even for a geodesic ball, under other curvature conditions.
Recently, the author [15] gave the first result on the Second Main Theorem in Nevanlinna theory on complete Kähler manifolds with non-negative Ricci curvature by introducing the global Green function method. Consequently, the author answered the Picard’s problem affirmatively provided that is of maximal volume growth:
where denotes the Riemannian volume of a geodesic ball centered at a fixed reference point with radius in
The original technique in [15] is to construct a family of relatively compact exhaustive domains of using the minimal positive global Green function for Based on an asymptotic estimate of obtained by Colding-Minicozzi [9], one can obtain two-sided bounds of the local Green function for in terms of integral forms. As a result, the author established the Second Main Theorem of meromorphic mappings on Let us introduce his main result as follows.
Colding-Minicozzi [9] (see [25] also) described the asymptotic behavior of : there exists a constant such that
where denotes the Riemannian distance function of from By means of the asymptotic estimate, the author defined a family of relatively compact domains as follows
for Hence, the Green function of for with a pole at satisfying Dirichlet boundary condition can be formulated as
According to Lemma 3.5 in [15], for any there exists such that
holds for all satisfying Based on this estimate, the Calculus Lemma was established. For is conformally equivalent to since has non-negative Ricci curvature and maximal volume growth (see [25]). So, by employing the standard arguments, the author established the Second Main Theorem of meromorphic mappings from into a complex projective manifold (see [15], Theorem 3.10), which gives an immediate result:
Theorem D (Dong, 2023).
If is of maximal volume growth, then every meromorphic function on must be a constant if it omits distinct values.
However, the technique in [15] depends on Colding-Minicozzi’s asymptotic estimate. Unfortunately, there are no such asymptotic estimates for if is not of maximal volume growth.
In this paper, we will generalize the Nevanlinna theory in the case that is a non-parabolic manifold, i.e., there exists a positive global Green function for Note that is non-parabolic if is of maximal volume growth. By N. Varopoulos [45, 46] (see Li-Tam-Wang [23, 25] also), the non-parabolicity implies that there exists a unique minimal positive global Green function of for denoted also by Our original idea is the construction of domain and application of Li-Yau’s estimate on heat kernel and global Green function. Using Li-Yau’s estimate [26], there exist constants such that
holds for all and For we redefine
Then, we have
which defines the Green function of for with a pole at satisfying Dirichlet boundary condition. There is a natural relationship between and : for we have
which plays an important role in the establishment of Nevanlinna theory in this paper (see Lemma 4.1).
Negative Curvature Method and Logarithmic Derivative Lemma Method are very typical approaches in the study of Nevanlinna theory. In this paper, the Logarithmic Derivative Lemma Method is used. So, we need to establish the so-called Calculus Lemma which is a way to the Logarithmic Derivative Lemma. We will introduce two techniques in the paper: Gradient Estimates (see Theorem 4.2) and Calculus Lemma (see Theorem 5.2) as follows.
Gradient Estimates. For any there exists such that
holds for all
The above gradient estimate yields an estimate of the harmonic measure on with respect to : for any there exists such that
holds for all (see Corollary 4.3).
Calculus Lemma. Let be a locally integrable function on Assume that is locally bounded at Then for any there exist a constant and a subset of finite Lebesgue measure such that
holds for all outside
Let be a nonconstant meromorphic function on The Calculus Lemma can derive the Logarithmic Derivative Lemma: for any there exists a subset of finite Lebesgue measure such that
holds for all outside (see Theorem 5.6).
@@ 1.2. Main results
Let be a complex projective manifold with over which one can put a Hermitian positive line bundle with Chern form We now take a divisor where denotes the complete linear system of Given a meromorphic mapping we have the Nevanlinna’s functions and (see Section 3.2). Let be the Ricci form of associated with the metric and be the canonical line bundle over The characteristic function of is defined by
The main result is the Second Main Theorem as follows.
@@
holds for all outside
@@
Since then we obtain a defect relation:
@@
Let denote the scalar curvature of Set
In the non-parabolic case, we answer the Picard’s problem as follows.
Corollary 1.3 (=Corollary 6.10).
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. Every meromorphic function on reduces to a constant if it omits distinct values, where denotes the maximal integer not greater than
It implies that every meromorphic function with on reduces to a constant if it omits 2 distinct values.
@@ 2.1. Volume Comparison Theorem
When Calabi-Yau (see [39]) gave a lower bound of
@@
2.2. Estimates of Heat Kernels
@@
where is the Dirac function with a pole at
When is non-parabolic with non-negative Ricci curvature, Li-Yau [26] (see [39] also) obtained an estimate of as follows.
@@ 3.1. Construction of
we see that is a bounded domain containing such that
Furthermore, the family exhausts i.e., for any sequence with we have
@@ 3.2. Nevanlinna’s Functions
Equip with the Kähler metric Let be the Hilbert-Schmidt norm of differential with respect to the metrics We have
@@ 4. Gradient Estimates of Green Functions
then we obtain
Let denote the gradient operator on Then, we have an upper bound of as follows.
@@
For any there exists such that
holds for all where is given by
@@
It yields from (1) that
Since
then the L’Hospital’rule yields that
where is the cut locus of and is the inward normal derivative on By
we see that for any there exists such that
holds for all Since has measure 0, with the aid of the continuity of on and on one then has the theorem proved. ∎
As a result, we obtain an estimate for upper bounds of harmonic measure on with respect to as follows:
@@
For any there exists such that
holds for all where is given by
@@ 5.1. Calculus Lemma
Note that exists for almost every Then, we have
@@
Let be a locally integrable function on Assume that is locally bounded at Then for any there exist a constant and a subset of finite Lebesgue measure such that
holds for all outside
@@
Apply Borel’s lemma to the left hand side of the above equality twice: one is to and the other is to we see that for any there exists a subset of finite Lebesgue measure such that
holds for all outside On the other hand, Corollary 4.3 implies that there exists such that
holds for all where is given by (1). Combining the above, we have
holds for all outside Therefore, we have the theorem proved by setting ∎
Corollary 5.3.
Let be a locally integrable function on Assume that is locally bounded at Assume that is of maximal volume growth. Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
@@ 5.2. Logarithmic Derivative Lemma
in a holomorphic local coordinate where is the inverse of Again, define the Nevanlinna’s characteristic function of by
@@
holds for all outside
@@
Next, we establish the following Logarithmic Derivative Lemma.
@@
holds for all outside
@@
holds for all outside ∎
By we have:
Corollary 5.7.
holds for all outside
@@ 6.1. Second Main Theorem
holds for all outside
@@
Since is of simple normal crossing type, then we can provide a finite open covering of and finitely many rational functions on for every such that are holomorphic on with that
Moreover, we may require for On write
@@
Fix any we can choose a holomorphic local coordinate near and a holomorphic local coordinate near such that
@@
for all outside By this with (6.1), we prove the theorem. ∎
Corollary 6.2.
holds for all outside
6.2. Defect Relation
The simple defect of with respect to is defined by
@@ 6.2. Defect Relation
Let
denote the Ricci curvature tensor of Since is Kählerian, then
Again, denote by the scalar curvature of Note that is the trace of thus we have
It yields that
Corollary 6.4.
Let be a meromorphic mapping. If there exists a number such that satisfies the growth condition
for sufficiently large then must be differentiably degenerate.
@@
Assume on the contrary that is differentiably non-degenerate. Then, we obtain as The growth condition implies that
Thus, we deduce due to Corollary 6.3, but which contradicts with ∎
Consider a compact Riemann surface of genus Note that
for an arbitrary point where is the holomorphic line bundle over defined by Combining this with Theorem 6.1, we conclude that
Corollary 6.5.
Let be a nonconstant meromorphic mapping. Let be distinct points in Then
Let
be a meromorphic mapping. Then
Let be the hyperplane line bundle over Since
then Theorem 6.1 yields that
Corollary 6.6.
Let be a differentiably non-degenerate meromorphic mapping. Let be hyperplanes in general position in Then
A meromorphic function on can be seen as a meromorphic mapping Set
We obtain a Picard’s theorem as follows.
Corollary 6.7 (Picard’s Theorem).
Every meromorphic function on must be a constant if it omits distinct values, where denotes the maximal integer not greater than
In the following, we shall derive a better Picard’s theorem compared with Corollary 6.7. On we introduce the spherical distance
Let be a meromorphic function on which is regarded as a meromorphic mapping
In terms of the spherical distance, the proximity function of with respect to can be defined as
Adopt the Ahlfors-Shimizu’s characteristic function of as follows
Define
Then, we have the First Main Theorem
Theorem 6.8 (Second Main Theorem).
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. Let be distinct values in Let be a nonconstant meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
Proof.
Define
Set
Clearly, we see that
Corollary 6.9 (Defect Relation).
Assume the same conditions as in Theorem Then
We derive a better Picard’s theorem as follows.
Corollary 6.10 (Picard’s Theorem).
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. Every meromorphic function on must be a constant if it omits distinct values, where denotes the maximal integer not greater than

### v7

# On Picard’s Problem via Nevanlinna Theory

@@ Abstract.
We consider the classical Picard’s problem for non-parabolic complete Kähler manifolds with non-negative Ricci curvature. Based on the global Green function approach, we give a positive answer to Picard’s problem under certain condition by developing Nevanlinna theory. That is, we prove that every meromorphic function on such a manifold reduces to a constant if it omits three distinct values, provided that the manifold satisfies a volume growth condition.
@@
Nevanlinna theory; value distribution; Picard’s theorem; defect relation; Kähler manifolds2010 Mathematics Subject Classification:
32H30; 32H25; 32A221. Introduction
1.1. Motivation
The famous Picard’s theorem asserts that a meromorphic function on the complex Euclidean spaces must reduce to a constant if it omits three distinct values. Historically, many authors have been committed to extensions of this theorem for a long time (see [4, 13, 15, 17, 18, 21, 28]). The study of Picard’s problem on a manifold may have originated from the celebrated work due to S. T. Yau [46] on Liouville’s theorem, which proves the Liouville’s property of harmonic functions on Ricci non-negatively curved manifolds. Yau’s work inspired people to think about the following Picard’s problem. Let be a complete non-compact Kähler manifold with non-negative Ricci curvature (see examples in Sha-Yang [35] and Tian-Yau [41, 42]). The classical Picard’s problem asks that
@@ 1.1. Motivation
Remark 1. Let be the universal covering, where is the unit disc in Thanks to the Liouville’s theorem by Yau, one may give an affirmative answer to the Picard’s problem if every holomorphic mapping can lift to a holomorphic mapping by Using the mapping lifting theorem, admits a lifting if and only if
for some such that Since is simply-connected, then the fundamental group is trivial. It implies that is necessarily a trivial group if such a lifting exists. However, is not trivial in most cases. So, we cannot lift to in general.
An early result in this direction could be dated back to 1970, S. Kobayashi [21] obtained some Picard-type theorems for holomorphic mappings between certain complex manifolds. In particular, he showed that
@@
Kobayashi’s result is subject to certain complex manifolds which are acted on transitively by a complex Lie group. In 1975, Goldberg-Ishihara-Petridis [18] treated a class of locally flat manifolds (see N. Petridis [28] also). They showed that
@@
If is locally flat, then every holomorphic mapping of bounded dilatation must be a constant.
It is known that Nevanlinna theory [27] is a great development of Picard’s theorem, which investigates the value distribution of meromorphic mappings between complex manifolds. Refer to L. Ahlfors [5], H. Cartan [7], Carlson-Griffths-King [8, 16], J. Noguchi [24, 25], E. Nochka [26], M. Ru [29, 30], B. Shiffman [31], B. Shabat [32], F. Sakai [33, 34], W. Stoll [39, 40], P. Vojta [38] and H. Wu [45], and see [1, 2, 3, 10, 11, 12] and references therein.
In 2010, A. Atsuji [2] gave an affirmative answer for a class of meromorphic functions of slow growth rate by establishing a Nevanlinna-type theory based on heat diffusion, We shall introduce his main work. Let be the Kähler form and Laplace-Beltrami operator of associated to respectively. One can endow with the Fubini-Study metric Let be a meromorphic function. The Kählerness of implies the Hilbert-Schmidt norm of the differential (with respect to ) can be formulated as
@@
Atsuji’s method [2] (see [10] also) is to use stochastic calculus by Brownian motion. Atsuji successfully introduced Nevanlinna-type functions and where is the time of the Brownian motion on By computing curvatures and utilizing Itô’s formula (see [19]), he established an analogue of Nevanlinna theory. Note that is stochastically complete (i.e., is conservative) since Grigor’yan’s criterion [14]. To make and meaningful, the following conditions are necessary:
@@
To do so, has to satisfy certain growth assumption. That is why a growth condition is needed in Theorem C.
In this paper, we generalize Nevanlinna theory to non-parabolic complete Kähler manifolds with non-negative Ricci curvature. Our innovative method is to construct a family of exhaustive precompact domains for by means of a unique minimal positive global Green function for so that we can well define the Nevanlinna’s functions on and establish a Second Main Theorem by introducing the global Green function method. Apply the Second Main Theorem, we can give a positive answer to the Picard’s problem under certain condition.
@@ 1.2. Main results
Assume that is non-parabolic, by which we mean that there exists a positive global Green function for By N. Varopoulos [43, 44] (see Li-Tam-Wang [20, 22] also), the non-parabolicity implies that there exists a unique minimal positive global Green function of for denoted by where is a fixed reference point in Using Li-Yau’s estimate [23], there are constants such that
holds for all and where is the Riemannian distance function of from and denotes the Riemannian volume of the geodesic ball centered at with radius in For define
Then, for all and the family exhausts
Let be a complex projective manifold with over which one can put a Hermitian positive line bundle with Chern form We now take a divisor where denotes the complete linear system of Given a meromorphic mapping we have the Nevanlinna’s functions and on (see Section 3.2). Let denote the Ricci form of associated to the metric and be the canonical line bundle over The characteristic function of is defined by
Again, set
The main result is the following Second Main Theorem.
@@
holds for all outside where is given by
@@
We obtain a defect relation:
@@
if one of the following conditions is satisfied
satisfies the volume growth condition
is of non-polynomial type growth, i.e., satisfies the growth condition
Remark 2. Note that due to
The condition (2) is relaxed. For example, it holds when with and holds when with or and
As an application of Corollary 1.2, we answer the Picard’s problem under certain condition as follows.
Corollary 1.3.
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. If satisfies the volume growth condition then every meromorphic function on reduces to a constant if it omits distinct values.
Corollary 1.4.
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. Then, every meromorphic function of non-polynomial type growth on can omit distinct values at most.
@@ 2.1. Volume Comparison Theorem
When Calabi-Yau (see [36]) proved that
@@
2.2. Estimates of Heat Kernel
@@
where is the Dirac’s delta function with a pole at
When is non-parabolic with non-negative Ricci curvature, Li-Yau [23] (see [36] also) gave two-sided bounds of as follows.
@@ 3.1. Construction of
where means we derive that is a precompact domain containing such that
In fact, the family exhausts namely, for any sequence with we have
@@ 4. Gradient Estimates of Green Functions
we obtain
Let denote the gradient operator on a Riemannian manifold. Cheng-Yau [9] proved the following theorem.
Lemma 4.2.
Let be a complete Riemannian manifold. Let be a non-negative harmonic function on a geodesic ball centered at with radius Then, there exists a constant depending only on the dimension of such that
where is the Riemannian distance between and and is the lower bound of the Ricci curvature on
We have an upper bound of as follows.
@@
There exists a constant such that
@@
By
we can obtain an estimate for upper bounds of as follows.
@@
There exists a constant such that
where is the Riemannian area element of
@@ 5.1. Calculus Lemma
Note that exists for almost all Then, we have
@@
Let be a locally integrable function on Assume that is locally bounded at Then there exists a constant such that for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is given by
@@
Apply Borel’s lemma to the left hand side of the above equality twice: one is to and the other is to then we deduce that for any there exists a subset of finite Lebesgue measure such that
holds for all outside On the other hand, Corollary 4.4 implies that there exists a constant such that
Set Combining the above, we have
holds for all outside ∎
@@ 5.2. Logarithmic Derivative Lemma
in holomorphic local coordinates here is the inverse of Again, define the Nevanlinna’s characteristic function of by
@@
holds for all outside where is given by
@@
We establish the following Logarithmic Derivative Lemma.
@@
holds for all outside where is given by
@@
holds for all outside where is given by ∎
@@ 6. Second Main Theorem and Defect Relation
holds for all outside where is given by
@@
Since is of simple normal crossing type, then there are a finite open covering of and finitely many rational functions on for each such that are holomorphic on with that
In addition, we can require that for all On write
@@
Fix any one can take holomorphic local coordinates near and holomorphic local coordinates near such that
@@
for all outside By this with (6), we prove the theorem. ∎
Recall that the simple defect of with respect to is defined by
@@
if one of the following conditions is satisfied
satisfies the volume growth condition
is of non-polynomial type growth, i.e., satisfies the growth condition
@@
Note that due to Corollary 2.2. We have
Using the volume growth condition, we obtain
Since a non-constant meromorphic mapping has the growth rate at least, then we have the theorem proved if is satisfied due to Theorem 6.1.
For we have
where
So, it yields from that
This completes the proof. ∎